OpenStudy (anonymous):

solve for x 1/6 (2x - 66) = 4x

OpenStudy (anonymous):

2x+6=66-4x Simple and best practice solution for 2x+6=66-4x equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for 2x+6=66-4x equation: Simplifying 2x + 6 = 66 + -4x Reorder the terms: 6 + 2x = 66 + -4x Solving 6 + 2x = 66 + -4x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right.

OpenStudy (anonymous):

@Swagger.. i think that you got the question mixed up its 1/6 ahead of all of the x variables

OpenStudy (anonymous):

1/3x+11=4x 11=4x-1/3x 11=11/3x x=3 not so sure

OpenStudy (anonymous):

@cassytaylor92 its just a negative but u got it right

OpenStudy (anonymous):

1/3x-11=4x -11=4x-1/3x -11=11/3x x=-3

OpenStudy (anonymous):

i'm sorry @Starr_Dynasty

OpenStudy (anonymous):

u did awesome @cassytaylor92